jigglypuff314 (jigglypuff314):

"Find the cube roots of 27(cos 330 + i sin 330)" Please help? I am completely clueless on what it is asking or how to go about solving it. (I'm asking this for a friend)

OpenStudy (alekos):

are you familiar with cosx + isinx = e^ix ?

OpenStudy (anonymous):

divide by 3

jigglypuff314 (jigglypuff314):

idk but this was a given http://prntscr.com/6213hi

OpenStudy (anonymous):

well also take the cube root of 27, which is 3 since \(3^3=27\) i meant "divide the angle by 3"

OpenStudy (anonymous):

that is a bunch of math teacherse designed to confuse something that is very very (very) easy

OpenStudy (anonymous):

divide the angle by 3, you can do it in your head

jigglypuff314 (jigglypuff314):

110 degrees?

OpenStudy (anonymous):

bingo

OpenStudy (alekos):

that's De'Moivres theorem which satellite is trying to explain to you

OpenStudy (anonymous):

so one answer is \[3(\cos(110^\circ)+i\sin(110^\circ))\]

OpenStudy (anonymous):

now you have a choice of ways to find the other two answers

OpenStudy (anonymous):

one is to divide the unit circle in two three equal parts with \(110^\circ\) as one of the parts

jigglypuff314 (jigglypuff314):

I have to show work, what is this De'Moivres theorem?

OpenStudy (anonymous):

the other is to add 360 and divide by 3 again

OpenStudy (anonymous):

show work: \(\sqrt[3]{27}=3\) and \(330\div 3=110\) is the first part of the "work"

jigglypuff314 (jigglypuff314):

haha ok :)

OpenStudy (anonymous):

ok now here is an interesting question before we continue

jigglypuff314 (jigglypuff314):

oh! the weird hieroglyphics are starting to make sense!

OpenStudy (anonymous):

most adults do this using radians (numbers) and not degrees in fact the formula you gave as a screen shot is done in radians but the question is asked in degrees

jigglypuff314 (jigglypuff314):

I realized that xD hehe what's the k in that equation?

OpenStudy (anonymous):

well degrees we have, so degrees we work with for the next bit of work, take \(330+360\) and divide it by \(3\)

jigglypuff314 (jigglypuff314):

330 +360 = 690 /3 = 230

OpenStudy (anonymous):

the \(k\) is \(0,1,2,3...\) however many you need for the roots we will use 0,1, 2

jigglypuff314 (jigglypuff314):

is that the new angles we put in there? would we still keep it as 3 on the outside part?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the real cube root of 27 does not change

jigglypuff314 (jigglypuff314):

so 3(cos(230) + isin(230)) ?

OpenStudy (anonymous):

bingo one more

jigglypuff314 (jigglypuff314):

add 360 again?

OpenStudy (anonymous):

yup

jigglypuff314 (jigglypuff314):

to the 230 or the 690?

OpenStudy (anonymous):

to the 690

OpenStudy (anonymous):

in other words \(330+2\times 360\) this time \(k=2\)

jigglypuff314 (jigglypuff314):

ok :) 690 + 360 = 1050 1050 / 3 = 350

OpenStudy (anonymous):

you found the sound

jigglypuff314 (jigglypuff314):

I think I'm getting the hang of this now! :D

OpenStudy (anonymous):

let me draw you a picture

jigglypuff314 (jigglypuff314):

so the final one is 3 (cos(350) + isin(350))

OpenStudy (anonymous):

yes

OpenStudy (anonymous):



OpenStudy (anonymous):

we divided the unit circle in to three parts, with one at \(110^\circ\)

OpenStudy (anonymous):

i should say "three equal parts"

jigglypuff314 (jigglypuff314):

unit circle divided into thirds to find the cubes?

OpenStudy (anonymous):

yes, so long as you know one of them to start you can do it that way

OpenStudy (anonymous):

we had to find the 110 first, but that was an easy matter of division

OpenStudy (anonymous):

no prized for guessing how many part for fourth roots etc

OpenStudy (anonymous):

*prizes

OpenStudy (anonymous):

ooh i should correct what you said divide in to three parts for cubed roots, not cubes

jigglypuff314 (jigglypuff314):

oh right ^_^

OpenStudy (anonymous):

now go show your friend how easy this was

jigglypuff314 (jigglypuff314):

I will! :D Thank you so much for your help!

OpenStudy (anonymous):

it is the whole reason to write a complex number in this form at least one main reason

OpenStudy (anonymous):

yw